an object moves in a straight line northward at a speed of 5m/sec .and its speed reaches 20m/sec . through 3seconds (a)the velocity after 3 seconds
1
Expert's answer
2021-10-21T16:33:23-0400
"a=\frac{20-5}{3}=5\frac{m}{s^2}\\V=20+5(3)=35\frac{m}{s}"